### v1

# Refining Noise Mitigation in NISQ Hardware Through Qubit Error Probability

Abstract
A key characteristic of the Noisy Intermediate-Scale Quantum (NISQ) era is the accurate assessment and consideration of errors through error mitigation techniques, which play a fundamental role in achieving quantum utility, i.e., quantum results that are at the state of the art of classical calculations. In this paper, we examine the primary sources of errors in contemporary (IBM) quantum computers through a new important metric, the qubit error probability (QEP), which estimates the probability of a qubit to suffer an error. We show that QEP can be used to improve one of the most important error mitigation techniques, the zero noise extrapolation (ZNE). Our method, named zero error probability extrapolation (ZEPE), is based on calibration parameters which enables a good scalability in terms of number of qubits and circuit depth. Besides, we prove that ZEPE performs better than the standard ZNE, specially for a mid-size depth ranges.
  • March 2025
1 Introduction
Quantum computing, defined as the ability to harness real quantum states to perform complex calculations, has recently transitioned from a theoretical possibility to a practical reality [1]. Feynman’s proposition [2] of using quantum systems to simulate quantum mechanics is no longer a conceptual vision but a tangible achievement. Over the past two decades, the capabilities of advanced quantum computers have seen remarkable progress. In principle, ideal quantum computers are expected to execute calculations beyond the reach of classical computers with exceptional precision. However, achieving this accuracy requires devices with thousands of qubits to enable effective quantum error correction [3, 4, 5]. Unfortunately, this goal remains distant. In the current noisy intermediate-scale quantum (NISQ) era, scientific and technological efforts are focused on evaluating, controlling, and mitigating physical errors [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
NISQ devices typically consist of a few dozen qubits and are expected to scale to a few hundred in the near future. A defining characteristic of these devices, as their name implies, is their inherent noisiness. Currently, the most advanced quantum computers are based on transmon qubits [23]. However, these systems face five significant limitations. First, the stability of a qubit state in these devices is typically of the order of hundreds of microseconds, thus restricting the duration of feasible computations. Second, the operations (gates) that act on the qubits are noisy, leading to a progressive loss of accuracy during calculations. Third, the process of measuring the qubit state is prone to errors. Fourth, there is a possibility that an operation on one qubit affects another, the crosstalk. Finally, the number of available physical qubits remains limited, imposing further constraints on computational capability.
Normally, when we want to estimate the error of any quantum calculation, we compare it with the simulated one. Nevertheless, quantum computers in the NISQ era are able to perform state-of-the-art classical computations [24, 25] and it is expected that they may overcome them in the near future. For this reason, it is preferable to avoid relying on classical computations to determine the error in quantum calculations. Therefore, we need to find metrics to measure errors that avoid the use of classical simulations. Sometimes we can use specific Clifford points [24] where the exact analytical result is known and then extrapolate the results to the non-trivial points. Nevertheless, there are many cases in which we may not be able to compare this with any analytical result. In a previous paper [26] we proposed a pre-processing tool (the Tool for Error Description in Quantum Circuits, TED-qc), which allows us to calculate the probability of an error to occur in a quantum circuit provided the hardware and its calibration, without any actual evaluation of the circuit. This tool has proven to be a good estimator of the error of a quantum circuit [27, 28, 29, 30]. However, when a calculation involves many qubits, its usefulness may be limited, as errors in the quantum circuit do not necessarily impact the final result in all cases. In this paper, we introduce a more precise metric for assessing the actual error in a quantum computation. Rather than considering only the total error probability, we focus on the individual qubit error probability (QEP), which provides a more refined measure of the error’s impact on the calculation.
As noted above, quantum computing faces significant challenges due to noise and errors that arise from imperfections in hardware, environmental interactions, and quantum decoherence. Unlike classical computers, which can use traditional error correction methods, quantum systems require specialized techniques to mitigate noise while preserving quantum coherence. Error mitigation techniques aim to improve the accuracy of quantum computations in NISQ devices at the cost of increasing the quantum computation time. This may allow researchers and industry professionals to extract meaningful results despite the presence of noise. These techniques do not eliminate noise entirely, but instead reduce its effects, making quantum algorithms more reliable and practical for real-world applications.
Several error mitigation strategies [31] have been developed to address different sources of errors, including Probabilistic Error Cancellation (PEC), which leverages classical post-processing to counteract noise by applying carefully designed inverse transformations, and Dynamical Decoupling, which use precisely timed control pulses to suppress decoherence and extend qubit coherence times. Additionally, measurement error mitigation techniques correct readout errors by characterizing and inverting the noise model affecting quantum measurements.
Among the error mitigation techniques, one of the most widely used is the Zero-Noise Extrapolation (ZNE) [32, 33, 34], which systematically amplifies noise in a controlled manner, executes the quantum circuit under varying noise regimes, and extrapolates the computed results to approximate an idealized zero-noise limit. This approach relies on the assumption that the quantum system’s response to noise follows a predictable mathematical trend, enabling the reconstruction of an estimated noiseless expectation value. The fundamental principle of ZNE is to employ numerical extrapolation methods —such as linear, polynomial, or exponential regression— to model the dependency of the circuit’s output on noise strength and infer the limit where noise approaches zero. By artificially scaling noise levels, typically through pulse stretching, gate repetition, or hardware-level noise injection, ZNE enables the extraction of noise-free estimates without requiring full quantum error correction. This technique has demonstrated effectiveness in improving the fidelity of quantum computations in domains such as quantum chemistry simulations, variational quantum eigensolvers, and other near-term quantum algorithms where mitigating errors is essential [24, 35, 36, 37]. Importantly, while the computational cost of techniques such as PEC increases exponentially with the number of qubits, ZNE remains independent of qubit count, typically requiring only a error amplification factor of 3-5 in additional quantum computational resources. This scalability makes it one of the most applied quantum error mitigation techniques.
The key issue regarding ZNE is how errors are quantified and amplified. The usual procedure, like the one implemented in frameworks such as Qiskit, is to consider that the original circuit has an error of one. Then, copies of the circuit are introduced in pairs, preserving the computed result while increasing the overall error. These methods use a metric that considers the error in a simplistic manner; for instance, if a new circuit is built with three copies of the original circuit, it is assumed that the error is three times the original. This is a dubious approximation since the error of a circuit does not increase linearly with the circuit depth as we will see later on. In this paper, we propose to use the mean qubit error probability (QEP) as a metric to quantify and control error amplification. This metric represents more accurately the error that occurs in quantum circuits.
Our benchmark circuit to test both the QEP and its effect in ZNE is the Trotterized time evolution of a two-dimensional transverse-field Ising model. The Ising model is widely studied across various domains of physics and has been extended in innovative ways in recent simulations investigating quantum many-body phenomena, including time crystals, quantum scars, and Majorana edge modes[38, 39, 40, 41].
Specifically, we examine the time dynamics of the Hamiltonian:
$H=-J\sum_{\left<i,j\right>}Z_{i}Z_{j}+h\sum_{i}X_{i},$ (1)
where $J>0$ is the coupling of nearest-neighbor spins with $i<j$ and $h$ is the global transverse field. Here, $Z_{i}$ and $X_{i}$ represent the Pauli-$Z$ and Pauli-$X$ operators acting on qubit $i$, respectively. The time evolution of an initial quantum state can be approximated using a first-order Trotter decomposition of the time-evolution operator, which factorizes the Hamiltonian dynamics into discrete steps, in the following way:
[…]
$e^{-iH_{X}\delta_{t}}$ $=$ $\prod_{i}e^{-ihX_{i}\delta_{t}}=\prod_{i}R_{X_{i}}(2h\delta_{t}).$ (2)
where $H_{ZZ}=-J\sum_{\left<i,j\right>}Z_{i}Z_{j}$ and $H_{X}=h\sum_{i}X_{i}$ denote the two separate terms of the Hamiltonian, which are independently exponentiated in the Trotter decomposition. The evolution time $T$ is discretized into $T/\delta t$ Trotter steps, and $R_{Z_{i}Z_{j}}(\theta_{j})$ and $R_{X_{i}}(\theta_{h})$ are ZZ and X rotation gates, respectively. Each Trotter step sequentially applies the two evolution operators defined above, and the total circuit depth increases proportionally with the number of Trotter steps. This flexibility allows for controlled simulation accuracy by increasing the number of steps.
We have chosen $\delta_{t}=1/4$, $J=\pi$, and $h=0$ to position the system at a Clifford point, which enables an efficient comparison between our simulations and the exact analytical solution. The chosen expectation value is the total magnetization $M=\sum_{i}\langle Z_{i}\rangle$.
The paper is organized as follows. In section 2, we introduce the new qubit error probability and compare it with other metrics. In section 3, we introduce a new extrapolation technique, the zero error probability extrapolation (ZEPE) method, to improve standard ZNE. In section 4, we compare the results obtained using our new tool against prior methods. Finally, in section 5, we summarize our conclusions.
2 Qubit error probability
As discussed in the previous section, error estimation is crucial in the NISQ era of quantum computing. In [26], we calculated the total error probability and, following the same scheme, we will calculate the qubit error probability (QEP). As we know, errors arise primarily from four sources:
  • The error induced by the instability of the qubits. In order to have non-trivial states we have to excite our physical qubits. The probability of finding the qubit in the excited state decays exponentially with time [42, 43], so the larger the quantum circuit the less likely is to find the qubit in the expected state. There are two possible decaying mechanisms, the decay of an excited state to the ground state, i.e. the probability of a state $|1\rangle$ to decay to a $|0\rangle$; and the change of the phase of an excited state, for example passing from the state $1/\sqrt{2}(|0\rangle+|1\rangle)$ to the state $1/\sqrt{2}(|0\rangle-|1\rangle)$.
  • Gate errors: each quantum gate introduces errors with single-qubit gates typically having error probabilities of the order of $10^{-4}$ to $10^{-3}$ and two-qubit gates of the order of $10^{-3}$ to $10^{-2}$.
  • Measurement errors: the act of measuring qubits introduces additional inaccuracies, typically on the order of $10^{-2}$ to $10^{-1}$.
  • The crosstalk between different qubits: when we act on a qubit we may act unintentionally on another. This type of error is very difficult to track, but luckily the latest IBM processor, Heron, have almost zero crosstalk [44, 45, 46]. All the results shown in this paper have been calculated on IBM Quantum Heron processor on the $ibm\_torino$ quantum system.
In order to calculate the qubit error probability we define the success probability of the j-th qubit, i.e. the possibility of non-committing any error on the j-th qubit, $S_{j}$, as
$S_{j}=\prod_{i=1}^{m}S_{ij}=\prod_{i=1}^{m}[1-P_{ij}],$ (3)
where $P_{ij}$ represents the error probability from any source acting on the j-th qubit: single and double qubit gates, measurement, and qubit instabilities; where $m$ is the total number of error sources. The relation between error probability and success probability is defined as $P+S=1$. The qubit error probability, QEP is therefore computed as
$P_{j}=1-[1-P_{j}^{\mathrm{meas}}][1-P_{j}^{\tau_{1}}][1-P_{j}^{\tau_{2}}]\prod_{i=1}^{\mathrm{gates}}[1-P_{ij}^{\mathrm{gate}}].$ (4)
The first term is the measurement error for the j-th qubit, denoted as $P_{j}^{\mathrm{meas}}$, and the second and third terms are the errors arising from qubit instability, represented by $P_{j}^{\tau_{1}}$ and $P_{j}^{\tau_{2}}$. Finally, the error due to all gates applied/affecting on the j-th qubit are considered on the last term, $P_{ij}^{\mathrm{gate}}$. While the gate and measurement errors are directly obtained through the IBM’S API, to account for errors caused by instability qubits $P_{j}^{\tau_{i}}$ we consider an exponential decay of the form [47]
$P_{j}^{\tau_{i}}=1-e^{-\frac{t_{j}}{\tau_{i,j}}}$ (5)
where $i=1,2$ are the two instability mechanism explained before, $t_{j}$ is the total circuit time of j-th qubit and $\tau_{i,j}$ are the relaxation and dephasing times of the j-th qubit. The evaluation of equation 5 requires the knowledge of the time it takes for any qubit between the initialization of the circuit and the measurement $t_{j}$. This is done by adding up how long it takes for the j-th qubit to perform any single gate. The situation becomes more complex when dealing with two-qubit gates. The process of executing a two-qubit gate necessitates the completion of all preceding gate operations on both qubits before the joint operation can be executed. Consequently, after incorporating the time needed for the two-qubit gate into the total time calculation, we are faced with the task of comparing the cumulative times of the two qubits involved. The crucial point here is that we don’t simply add these times together. Instead, we analyze the total times for each qubit and then adopt the longer of the two as the updated qubit time. This approach ensures that we are accounting for the maximum time it could take, thus capturing the full potential for instability-induced error.
The last product in equation 4 accounts for all gate-related errors $P_{ij}^{\mathrm{gate}}$ (both single and double) that affect the j-th qubit. There are two different gates that may produce an error on the j-th qubit. On the one hand, we have the gates acting directly on it. On the other hand, we have the possibility that the control qubit on a two-qubit gate suffered an error while acting on it. In order to take this into account, we add the gates affecting the control qubit gate to the ones affecting the j-th one.
These errors depend on the specific quantum hardware, qubits and connections used. To address this, we developed TED-qc (Tool for Error Description in Quantum Circuits), an open-source pre-processing tool available on GitLab (https://gitlab.com/qjornet/ted-qc/). Currenlty written for Qiskit language, it analyzes quantum circuits, extracts error probabilities from IBM’s API and computes total error probabilities. The code can be easily extended to any quantum computer using its calibration parameters.
The calibration procedure of a quantum hardware can detect failed qubits and gates, and sometimes can return inaccurate values. We have included a warning in the code to be aware of the possibility of a gate being poorly calibrated. When a two-qubit gate has an error that is $100\%$ larger than the mean value of all the two-qubit gates involved in the circuit, we highlight this issue as poorly calibrated, since it can induce a huge qubit error probability. Sometimes there is no information regarding a particular gate, for example between the $85$th and $86$th qubits in figure 1a, hence QEP gives an error probability of $1$, which obviously may be misleading.
Figure 1: Schematic representation of the procedure to get the mean Qubit Error Probability. a) 12 phyiscal qubit mapping on $ibm\_torino$ quantum system for the 12 site Ising Hamiltonian (equations 1 and 2). b) to e) Computed Qubit error probability (QEP) of each qubit in a circuit with $1$, $5$, $10$, and $15$ Trotter steps, respectively. Dashed red and green lines represent the mean and standard deviation of the QEP values, respectively.
Figure 1 shows an example of how QEP values are obtained in our tool. We have run a $12$-qubit trotterization of Ising Hamiltonian expressed in equations 1 and 2 for different numbers of Trotter steps on $ibm\_torino$. In the left panel a), we can see the 12-qubit chain chosen by the Qiskit transpile function required to map the logical qubits into the physical quantum hardware. In the right panels, b) to e), we can see the error probability of each qubit in the circuit for $1$, $5$, $10$, and $15$ Trotter steps, respectively. We can see how the mean and variance of the QEP increase as the number of Trotter steps are increased. It is also clear from the image that different qubits have different QEP; however, since the variance is one order of magnitude smaller in all the cases, we take the average value as the extrapolation variable for the error mitigation strategy.
3 Using Qubit Error Probability in the Zero Noise Extrapolation, the ZEPE method
Once we have defined the QEP, we can use it as the extrapolation variable in the ZNE method. We refer to this new ZNE extrapolation method as Zero Error Probability Extrapolation (ZEPE). The idea is the same as in ZNE: we execute multiple circuits with the same expectation value as the original one but ensuring that the replicas will have larger error probabilities each time. To do so, we insert pairs of two-qubit gates on all connected qubits in the circuit. We know that a pair of two-qubit gate is just a unitary operation, so the expectation value remains the same as the original one. The addition of this set of gates defines a factor of one. We can add as many factors as we want to further amplify the noise, and hence the error probability (see figure 2). As we pointed out before, we use the mean QEP to measure the ”error”, and then we extrapolate the mean QEP to zero in order to correct our result.
Figure 2: Transpiled quantum circuit for the Ising Hamiltonian evolution (equations 1 and 2) for a 8 qubits system. In blue are represented the basic gate operations used in $ibm\_torino$ quantum system for two (dark and light blue) trotterization steps. In red, the CZ entangle gates used for the error amplification are represented. The different color intensity accounts for each of the two error amplification factor.
It is important to note that, as in ZNE, ZEPE is compatible with other error mitigation techniques. Here, we have used the T-REX mitigation technique [48, 49, 50] as implemented in Qiskit to eliminate the measurement errors. This is particularly relevant, because when we use ZEPE we cannot increase the error due to the measurement because the final state would not be invariant, meaning it cannot be mitigated within this framework. Since we have mitigated this source of error by applying T-REX, we did not take it into account while calculating the QEP, assuming it has been largely reduced.
It has been shown that ZNE performs better when combined with twirling techniques [51]. Since we aim to compare our method against ZNE, we have also incorporated this error mitigation technique in our calculations. Another important issue that we have to take into account is the possible calibration problems explained above. Then, the role of warnings identifying if a qubit suffers from calibration issues or unexpected behavior is crucial, since the QEP may be miscalculated, leading to ineffective mitigation.
Figure 3: Use of ZEPE to mitigate the error in calculating the magnetization of a 68 qubit system after $11$ Trotter steps in $ibm\_torino$. The green points are the raw, factor $1$, factor $2$, factor $3$ values of the magnetization as a function of the mean qubit error probability (QEP). The blue star represents the theoretical exact result, and the dashed red line is the linear error extrapolation. Finally, the green star stands for the mitigated expectation value obtained from the extrapolation to the zero-error probability.
In figure 3, we calculate the magnetization of an $11$-Trotter-step Ising system as a function of the average QEP. The first green point represents the result of the circuit without ZNE or ZEPE mitigation techniques. The other three green points correspond to the results obtained after adding a set of two-qubit gates as indicated above (factor $1$, factor $2$, and factor $3$). The blue star is the theoretical exact value, and the dashed red line represents the zero error probability extrapolation. We can see that this method effectively mitigates noise mostly due to the gate errors and qubit state decay. It is important to stress that this method, like ZNE, is heuristic, so we cannot guarantee that it will be able to reduce all the noise.
Figure 4: ZNE and ZEPE magnetizations as a function of the number of Trotter steps. Different ZEPE colors correspond to different extrapolation adjustments, i.e. using different set of error amplification factors. For instance, blue line labeled as (1,2,3,4,5) corresponds to the linear extrapolation performed with the raw, factor 1, factor 2, factor 3, factor 4, and factor 5 points.
Finally, we have analyzed ZEPE as a function of the error amplification factor. Different factors can be used to perform the extrapolation. We have run several quantum circuits increasing the number of Trotter steps using amplification error factor ranging from 1 to 5. Then, we used different combination of the expectation values of the amplified circuits to do the error zero probability extrapolation. Figure 4 shows the results of the extrapolated values using different error amplification factor for the QEP calculation. We can see that the best mitigation corresponds to the use of ”raw” and $1$, $2$, and $3$ factors to extrapolate.
Among all the extrapolation options, we have chosen to use the linear extrapolation for two reasons. First, we may expect that the error while calculating most operators, especially the magnetization, scales linearly with the mean QEP. Second, we aim to avoid extrapolations such as the exponential one, which can induce to a significant deviation from the actual expectation value.
4 The comparison between ZNE and ZEPE
Once we have developed the ZEPE method, we aim to compare it with the ZNE function implemented in Qiskit. As we have previously done, we compute a 68-qubit chain in $ibm\_torino$ quantum system for different numbers of Trotter steps. ZEPE has two main advantages in comparison with ZNE. First, using ZEPE, we have enhanced control over error and noise, utilizing the quantum computer’s calibration to more precisely monitor noise behavior. Second, increasing the noise is more straightforward and affordable, as it can be achieved by adjusting the circuit depth which enable us to apply it for more complex circuits.
Figure 5: Comparison on the reliability between the different error mitigation techniques. Left: Absolute value of the total magnetization as a function of the amount of Trotter steps. The raw (orange), ZNE (blue), ZEPE (green) and exact value (solid line at $M=68$) are included. Right: Relative error of the raw and extrapolated values compaired with the exact theoretical result.
In figure 5, we can see the absolute magnetization (left) of the raw, ZNE, and ZEPE results, and the absolute error (right) as a function of the number of Trotter steps. We can see that our method reduces the error more effectively than ZNE. ZEPE performs better for $6$-$16$ Trotter steps. The effectiveness of the method directly depends on the entangle depth of the original quantum circuit. On the one hand, if the raw circuit has a small error probability, i.e. short depth, we cannot mitigate a significant portion of it, since the currently NISQ quantum machine is accurate enough to handle it. On the other hand, if the raw circuit has a mean QEP larger than $0.6$, the error due to circuit time and the error propagation on the gate operations starts to be sufficiently large for an effective mitigation. In this regime, ZEPE performs better than ZNE because the latter suffers of incommensurate circuit time execution which yields in a inaccurate manner to account for noise.
We have shown that ZEPE can be an easily tuneable error mitigation technique. It requires minimal quantum computational resources and performs better than standard ZNE. This new method has been used on IBM quantum computers but can be extended to any type of hardware.
5 Conclusions
In this work, we have investigated the primary sources of errors in quantum computers to define the QEP, a fundamental metric for quantifying the likelihood of qubit errors. Given the inherent noise and instability in NISQ devices, to accurately estimate and mitigate these errors is essential for improving quantum computation reliability. The QEP metric provides a systematic approach to assessing errors at the qubit level, offering valuable insights into error propagation and its impact on quantum circuits. By leveraging QEP, we aim to enhance the precision and effectiveness of error mitigation strategies in quantum computing.
One of the most critical aspects of the NISQ era is the development of efficient error mitigation techniques to bridge the gap between noisy quantum hardware and the desired quantum advantage. In this context, we explored how QEP can be utilized to improve one of the most commonly used techniques, the zero-noise extraploation ZNE. Standard ZNE methods attempt to infer error-free results by executing quantum circuits at different noise levels and extrapolating back to a noiseless regime. However, these approaches have limitations, particularly in their sensitivity to noise models and the complexity of implementation. By integrating QEP into this process, we introduced a novel extrapolation method, ZEPE, which refines the standard ZNE approach and enhances its overall performance.
The results of our study demonstrate that ZEPE offers significant advantages over traditional ZNE methods. It requires minimal quantum computational resources while achieving higher accuracy in error mitigation. The adaptability and efficiency of ZEPE makes it a promising technique for improving quantum computations in the NISQ era. Additionally, ZEPE is easily tunable, allowing users to optimize its parameters based on the specific noise characteristics of a given quantum system. These attributes make it particularly valuable for practical implementations where computational overhead must be minimized while maintaining high levels of precision.
Although our research has primarily focused on IBM quantum computers, the ZEPE method is not hardware-specific and can be applied to other quantum computing platforms. This flexibility ensures that ZEPE can contribute to a wide range of quantum technologies, enhancing their error resilience and overall performance. Future work could explore further refinements to the technique, including its integration with other error mitigation methods and its application to more complex quantum algorithms. By advancing robust and efficient error mitigation strategies, such as ZEPE, we move closer to achieving practical quantum utility, where quantum computations can rival and surpass classical methods in real-world applications.
Acknowledgments
We acknowledge support from the Basq initiative. This work has been possible thanks to the quantum computational resources provided by the BasQ Strategy under the collaboration agreement between Ikerbasque Foundation and the University of the Basque Country (UPV/EHU) and the University on Mondragon, on behalf of the Department of Science, Universities and Innovation of the Basque Government. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team. We also acknowledge support from the Basque Goverment ELKARTEK KUBIT project with reference KK-2024/00105. N.S. acknowledges funding from the European Union under the Horizon Europe research and innovation programme (Marie Skłodowska-Curie grant agreement no. 101148213, EATTS).

### v2

@@
-Refining Noise Mitigation in NISQ Hardware Through Qubit Error Probability
+A Useful Metric for the NISQ Era: Qubit Error Probability and Its Role in Zero Noise Extrapolation
 Abstract
-A key characteristic of the Noisy Intermediate-Scale Quantum (NISQ) era is the accurate assessment and consideration of errors through error mitigation techniques, which play a fundamental role in achieving quantum utility, i.e., quantum results that are at the state of the art of classical calculations.
-In this paper, we examine the primary sources of errors in contemporary (IBM) quantum computers through a new important metric, the qubit error probability (QEP), which estimates the probability of a qubit to suffer an error.
-We show that QEP can be used to improve one of the most important error mitigation techniques, the zero noise extrapolation (ZNE).
-Our method, named zero error probability extrapolation (ZEPE), is based on calibration parameters which enables a good scalability in terms of number of qubits and circuit depth.
-Besides, we prove that ZEPE performs better than the standard ZNE, specially for a mid-size depth ranges.
-  • March 2025
-1 Introduction
-Quantum computing, defined as the ability to harness real quantum states to perform complex calculations, has recently transitioned from a theoretical possibility to a practical reality [1].
-Feynman’s proposition [2] of using quantum systems to simulate quantum mechanics is no longer a conceptual vision but a tangible achievement.
+Accurate assessment and management of errors is indispensable for extracting useful results from noisy intermediate-scale quantum (NISQ) devices.
+In this work, we propose the qubit error probability (QEP), a device specific metric that combines relaxation, dephasing, gate, and measurement contributions into a single per qubit figure of merit computable before execution.
+Leveraging QEP as the control variable, we revisit zero noise extrapolation (ZNE) by adding pairs of controlled native two-qubit gates on all connected qubit pairs to generate circuits with successively larger mean QEP; the zero error limit is then approximated by a linear regression of the measured observable against those values.
+Benchmarking on IBM Quantum Heron processors, we apply QEP guided ZNE to first order Trotterized simulations of the two dimensional transverse field Ising model, chosen as a representative interacting many body system, involving up to 68 qubits and 15 Trotter steps.
+In regimes where the raw circuits exhibit a finite mean QEP, the method suppresses observable errors beyond those attainable with circuit depth scaled ZNE, while requiring only three noise scaled evaluations and no additional classical post processing.
+These results demonstrate that QEP serves as a transparent and efficient error metric, and that its integration into ZNE provides a practical route to reliability gains on current superconducting hardware, without the resource costs associated with full quantum error correction.
+preprint: AIP/123-QED
+I Introduction
+Quantum computing, defined as the ability to harness real quantum states to perform complex calculations, has recently transitioned from a theoretical possibility to a practical reality Nielsen and Chuang (2011).
+Feynman’s proposition Feynman (1982) of using quantum systems to simulate quantum mechanics is no longer a conceptual vision but a tangible achievement.
 Over the past two decades, the capabilities of advanced quantum computers have seen remarkable progress.
 In principle, ideal quantum computers are expected to execute calculations beyond the reach of classical computers with exceptional precision.
-However, achieving this accuracy requires devices with thousands of qubits to enable effective quantum error correction [3, 4, 5].
+However, achieving this accuracy requires devices with thousands of qubits to enable effective quantum error correction Lidar and Brun (2013); Terhal (2015); Wendin (2017).
 Unfortunately, this goal remains distant.
-In the current noisy intermediate-scale quantum (NISQ) era, scientific and technological efforts are focused on evaluating, controlling, and mitigating physical errors [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22].
+In the current noisy intermediate-scale quantum (NISQ) era, scientific and technological efforts are focused on evaluating, controlling, and mitigating physical errors Bharti et al.
+(2022); Leymann and Barzen (2020); Porter and Joseph (2022); Kandala et al.
+(2017); Aspuru-Guzik et al.
+(2005); Cerezo et al.
+(2021); Arute et al.
+(2019); Preskill (2018); Xiao, Freericks, and Kemper (2021); Dalzell et al.
+(2020); Georgopoulos, Emary, and Zuliani (2021); Patel et al.
+(2020); Nation et al.
+(2021); Weidenfeller et al.
+(2022); Setiawan et al.
+(2021); Wu et al.
+(2021); Headley et al.
+(2022).
 NISQ devices typically consist of a few dozen qubits and are expected to scale to a few hundred in the near future.
 A defining characteristic of these devices, as their name implies, is their inherent noisiness.
-Currently, the most advanced quantum computers are based on transmon qubits [23].
+Currently, some of the most advanced quantum computers are based on transmon qubits Koch et al.
+(2007); Kim et al.
+(2023); AI and Collaborators (2024).
 However, these systems face five significant limitations.
@@
 Finally, the number of available physical qubits remains limited, imposing further constraints on computational capability.
-Normally, when we want to estimate the error of any quantum calculation, we compare it with the simulated one.
-Nevertheless, quantum computers in the NISQ era are able to perform state-of-the-art classical computations [24, 25] and it is expected that they may overcome them in the near future.
+The error estimation in a quantum calculation, usually requires comparing the result with the corresponding from a classical calculation.
+Nevertheless, quantum computers in the NISQ era are able to perform state-of-the-art classical computations Kim et al.
+(2023); AI and Collaborators (2024) and it is expected that they may overcome them in the near future.
 For this reason, it is preferable to avoid relying on classical computations to determine the error in quantum calculations.
 Therefore, we need to find metrics to measure errors that avoid the use of classical simulations.
-Sometimes we can use specific Clifford points [24] where the exact analytical result is known and then extrapolate the results to the non-trivial points.
-Nevertheless, there are many cases in which we may not be able to compare this with any analytical result.
-In a previous paper [26] we proposed a pre-processing tool (the Tool for Error Description in Quantum Circuits, TED-qc), which allows us to calculate the probability of an error to occur in a quantum circuit provided the hardware and its calibration, without any actual evaluation of the circuit.
-This tool has proven to be a good estimator of the error of a quantum circuit [27, 28, 29, 30].
+Sometimes we can use specific Clifford points Kim et al.
+(2023) where the exact analytical result is known and then extrapolate the results to the non-trivial points.
+However, there are many situations in which it is not possible to compare this with any analytical result.
+In a previous work Aseguinolaza et al.
+(2024) we proposed a pre-processing tool (the Tool for Error Description in Quantum Circuits, TED-qc), which computes the probability of an error to occur in a quantum circuit provided the hardware and its calibration, without any actual evaluation of the circuit.
+This tool has proven to be a good estimator of the error of a quantum circuit Brody and Sykes (2025); Gardeazabal-Gutierrez, Terres-Escudero, and Bringas (2024); Antero et al.
+(2025); de Keijzer et al.
+(2025).
 However, when a calculation involves many qubits, its usefulness may be limited, as errors in the quantum circuit do not necessarily impact the final result in all cases.
@@
 These techniques do not eliminate noise entirely, but instead reduce its effects, making quantum algorithms more reliable and practical for real-world applications.
-Several error mitigation strategies [31] have been developed to address different sources of errors, including Probabilistic Error Cancellation (PEC), which leverages classical post-processing to counteract noise by applying carefully designed inverse transformations, and Dynamical Decoupling, which use precisely timed control pulses to suppress decoherence and extend qubit coherence times.
+Several error mitigation strategies Cai et al.
+(2023) have been developed to address different sources of error, including Probabilistic Error Cancellation (PEC) Kim et al.
+(2023), which leverages classical post-processing to counteract noise by applying carefully designed inverse transformations, and Dynamical Decoupling, which use precisely timed control pulses to suppress decoherence and extend qubit coherence times.
 Additionally, measurement error mitigation techniques correct readout errors by characterizing and inverting the noise model affecting quantum measurements.
-Among the error mitigation techniques, one of the most widely used is the Zero-Noise Extrapolation (ZNE) [32, 33, 34], which systematically amplifies noise in a controlled manner, executes the quantum circuit under varying noise regimes, and extrapolates the computed results to approximate an idealized zero-noise limit.
+Among the error mitigation techniques, one of the most widely used is the Zero-Noise Extrapolation (ZNE) Li and Benjamin (2017); Temme, Bravyi, and Gambetta (2017); Giurgica-Tiron et al.
+(2020), which systematically amplifies noise in a controlled manner, executes the quantum circuit under varying noise regimes, and extrapolates the computed results to approximate an idealized zero-noise limit.
 This approach relies on the assumption that the quantum system’s response to noise follows a predictable mathematical trend, enabling the reconstruction of an estimated noiseless expectation value.
@@
 By artificially scaling noise levels, typically through pulse stretching, gate repetition, or hardware-level noise injection, ZNE enables the extraction of noise-free estimates without requiring full quantum error correction.
-This technique has demonstrated effectiveness in improving the fidelity of quantum computations in domains such as quantum chemistry simulations, variational quantum eigensolvers, and other near-term quantum algorithms where mitigating errors is essential [24, 35, 36, 37].
-Importantly, while the computational cost of techniques such as PEC increases exponentially with the number of qubits, ZNE remains independent of qubit count, typically requiring only a error amplification factor of 3-5 in additional quantum computational resources.
+This technique has demonstrated effectiveness in improving the fidelity of quantum computations in domains such as quantum chemistry simulations, variational quantum eigensolvers, and other near-term quantum algorithms where mitigating errors is essential Kim et al.
+(2023); Uvarov et al.
+(2024); Garmon, Pooser, and Dumitrescu (2020); Keen et al.
+(2020); Cai (2021a); He et al.
+(2020).
+Importantly, while the computational cost of techniques such as PEC or Richardson extrapolation Krebsbach, Trauzettel, and Calzona (2022); Russo and Mari (2024) increases exponentially with the number of qubits, ZNE remains independent of qubit count, typically requiring only a error amplification factor of 3-5 in additional quantum computational resources.
 This scalability makes it one of the most applied quantum error mitigation techniques.
@@
 Our benchmark circuit to test both the QEP and its effect in ZNE is the Trotterized time evolution of a two-dimensional transverse-field Ising model.
-The Ising model is widely studied across various domains of physics and has been extended in innovative ways in recent simulations investigating quantum many-body phenomena, including time crystals, quantum scars, and Majorana edge modes[38, 39, 40, 41].
-Specifically, we examine the time dynamics of the Hamiltonian:
+The Ising model is widely studied across various domains of physics and has been extended in innovative ways in recent simulations investigating quantum many-body phenomena, including time crystals, quantum scars, and Majorana edge modesCervera-Lierta (2018); Mi et al.
+(2022); Frey and Rachel (2022); Chen et al.
+(2022).
+Specifically, we examine the time evolution of the Hamiltonian:
 $H=-J\sum_{\left<i,j\right>}Z_{i}Z_{j}+h\sum_{i}X_{i},$ (1)
@@
 where $H_{ZZ}=-J\sum_{\left<i,j\right>}Z_{i}Z_{j}$ and $H_{X}=h\sum_{i}X_{i}$ denote the two separate terms of the Hamiltonian, which are independently exponentiated in the Trotter decomposition.
-The evolution time $T$ is discretized into $T/\delta t$ Trotter steps, and $R_{Z_{i}Z_{j}}(\theta_{j})$ and $R_{X_{i}}(\theta_{h})$ are ZZ and X rotation gates, respectively.
+The evolution time $T$ is discretized into $T/\delta t$ Trotter steps, and $R_{Z_{i}Z_{j}}(\theta_{j})$ and $R_{X_{i}}(\theta_{h})$ are the ZZ and X rotation gates, respectively.
 Each Trotter step sequentially applies the two evolution operators defined above, and the total circuit depth increases proportionally with the number of Trotter steps.
 This flexibility allows for controlled simulation accuracy by increasing the number of steps.
-We have chosen $\delta_{t}=1/4$, $J=\pi$, and $h=0$ to position the system at a Clifford point, which enables an efficient comparison between our simulations and the exact analytical solution.
-The chosen expectation value is the total magnetization $M=\sum_{i}\langle Z_{i}\rangle$.
+Following the lines of Kim Y.
+et al.Kim et al.
+(2023) we study the the total magnetization $M=\sum_{i}\langle Z_{i}\rangle$ for two different setup configurations.
+In the first one, we fix $\delta_{t}=1/4$, $J=\pi$, and $h=0$ to position the system at a Clifford point, allowing for an efficient comparison between our simulations and the exact analytical solution.
+In this case, we perform calculations up to 68 qubits and vary the number of Trotter steps.
+In the second one, we fixed the number of Trotter steps to $15$, $\delta_{t}=1/4$, $J=\pi$ for different values of $h$.
+As we are not at a Clifford point we have done the calculation up to $32$ qubits in order to be able to simulate the theoretical values in a classical computer.
 The paper is organized as follows.
-In section 2, we introduce the new qubit error probability and compare it with other metrics.
-In section 3, we introduce a new extrapolation technique, the zero error probability extrapolation (ZEPE) method, to improve standard ZNE.
-In section 4, we compare the results obtained using our new tool against prior methods.
-Finally, in section 5, we summarize our conclusions.
-2 Qubit error probability
+In section II, we introduce the qubit error probability (QEP), and compare it with other metrics.
+In section III, we use this QEP to track the error in order to improve standard ZNE.
+In section IV, we compare the results obtained using our new tool against prior methods.
+Finally, in section V, we summarize our conclusions.
+II Qubit error probability
 As discussed in the previous section, error estimation is crucial in the NISQ era of quantum computing.
-In [26], we calculated the total error probability and, following the same scheme, we will calculate the qubit error probability (QEP).
+In Aseginolaza U.
+et al.Aseguinolaza et al.
+(2024), we calculated the total error probability and, following the same scheme, we will calculate the qubit error probability (QEP).
 As we know, errors arise primarily from four sources:
@@
 In order to have non-trivial states we have to excite our physical qubits.
-The probability of finding the qubit in the excited state decays exponentially with time [42, 43], so the larger the quantum circuit the less likely is to find the qubit in the expected state.
+The probability of finding the qubit in the excited state decays exponentially with time Rigetti et al.
+(2012); Place et al.
+(2021), so the larger the quantum circuit the less likely is to find the qubit in the expected state.
 There are two possible decaying mechanisms, the decay of an excited state to the ground state, i.e.
@@
   • The crosstalk between different qubits: when we act on a qubit we may act unintentionally on another.
-This type of error is very difficult to track, but luckily the latest IBM processor, Heron, have almost zero crosstalk [44, 45, 46].
+This type of error is very difficult to track, but luckily the latest IBM processor, Heron, have almost zero crosstalk Sheldon et al.
+(2016); Dial (2022); Chow, Dial, and Gambetta (2021).
 All the results shown in this paper have been calculated on IBM Quantum Heron processor on the $ibm\_torino$ quantum system.
@@
 Finally, the error due to all gates applied/affecting on the j-th qubit are considered on the last term, $P_{ij}^{\mathrm{gate}}$.
-While the gate and measurement errors are directly obtained through the IBM’S API, to account for errors caused by instability qubits $P_{j}^{\tau_{i}}$ we consider an exponential decay of the form [47]
+While the gate and measurement errors are directly obtained through the IBM’s API, to account for errors caused by instability qubits $P_{j}^{\tau_{i}}$ we consider an exponential decay of the form Aleksandrowicz et al.
+(2019)
 $P_{j}^{\tau_{i}}=1-e^{-\frac{t_{j}}{\tau_{i,j}}}$ (5)
-where $i=1,2$ are the two instability mechanism explained before, $t_{j}$ is the total circuit time of j-th qubit and $\tau_{i,j}$ are the relaxation and dephasing times of the j-th qubit.
+where $i=1,2$ stand for the two instability mechanism, relaxation and dephasing explained before, $\tau_{i,j}$ are the related times for the j-th qubit, and $t_{j}$ is the total circuit time of j-th qubit.
 The evaluation of equation 5 requires the knowledge of the time it takes for any qubit between the initialization of the circuit and the measurement $t_{j}$.
@@
 To address this, we developed TED-qc (Tool for Error Description in Quantum Circuits), an open-source pre-processing tool available on GitLab (https://gitlab.com/qjornet/ted-qc/).
-Currenlty written for Qiskit language, it analyzes quantum circuits, extracts error probabilities from IBM’s API and computes total error probabilities.
+Currenlty written for Qiskit v1.2.0 language, it analyzes quantum circuits, extracts error probabilities from IBM’s API and computes total error probabilities.
 The code can be easily extended to any quantum computer using its calibration parameters.
@@
 Figure 1: Schematic representation of the procedure to get the mean Qubit Error Probability.
-a) 12 phyiscal qubit mapping on $ibm\_torino$ quantum system for the 12 site Ising Hamiltonian (equations 1 and 2).
+a) 12 phyiscal qubit mapping on $ibm\_torino$ quantum system for the 12 site Ising Hamiltonian (equation 1).
 b) to e) Computed Qubit error probability (QEP) of each qubit in a circuit with $1$, $5$, $10$, and $15$ Trotter steps, respectively.
-Dashed red and green lines represent the mean and standard deviation of the QEP values, respectively.
+Dashed red and green lines represent the mean ($\mu$) and standard deviation ($\sigma$) of the QEP values, respectively.
 Figure 1 shows an example of how QEP values are obtained in our tool.
-We have run a $12$-qubit trotterization of Ising Hamiltonian expressed in equations 1 and 2 for different numbers of Trotter steps on $ibm\_torino$.
+We have run a $12$-qubit trotterization of Ising Hamiltonian expressed in equation 1 with $\delta_{t}=1/4$, $J=\pi$, and $h=0$ for different number of Trotter steps on $ibm\_torino$.
 In the left panel a), we can see the 12-qubit chain chosen by the Qiskit transpile function required to map the logical qubits into the physical quantum hardware.
 In the right panels, b) to e), we can see the error probability of each qubit in the circuit for $1$, $5$, $10$, and $15$ Trotter steps, respectively.
-We can see how the mean and variance of the QEP increase as the number of Trotter steps are increased.
+We can see how the mean ($\mu$) and variance ($\sigma$) of the QEP increase as the number of Trotter steps are increased.
 It is also clear from the image that different qubits have different QEP; however, since the variance is one order of magnitude smaller in all the cases, we take the average value as the extrapolation variable for the error mitigation strategy.
-3 Using Qubit Error Probability in the Zero Noise Extrapolation, the ZEPE method
-Once we have defined the QEP, we can use it as the extrapolation variable in the ZNE method.
-We refer to this new ZNE extrapolation method as Zero Error Probability Extrapolation (ZEPE).
-The idea is the same as in ZNE: we execute multiple circuits with the same expectation value as the original one but ensuring that the replicas will have larger error probabilities each time.
-To do so, we insert pairs of two-qubit gates on all connected qubits in the circuit.
-We know that a pair of two-qubit gate is just a unitary operation, so the expectation value remains the same as the original one.
-The addition of this set of gates defines a factor of one.
-We can add as many factors as we want to further amplify the noise, and hence the error probability (see figure 2).
-As we pointed out before, we use the mean QEP to measure the ”error”, and then we extrapolate the mean QEP to zero in order to correct our result.
-Figure 2: Transpiled quantum circuit for the Ising Hamiltonian evolution (equations 1 and 2) for a 8 qubits system.
+III Using Qubit Error Probability in the Zero Noise Extrapolation
+Figure 2: Transpiled quantum circuit for the Ising Hamiltonian evolution (equation 1) for a 8 qubits system in the Clifford point ($J=\pi$, and $h=0$).
 In blue are represented the basic gate operations used in $ibm\_torino$ quantum system for two (dark and light blue) trotterization steps.
@@
 The different color intensity accounts for each of the two error amplification factor.
-It is important to note that, as in ZNE, ZEPE is compatible with other error mitigation techniques.
-Here, we have used the T-REX mitigation technique [48, 49, 50] as implemented in Qiskit to eliminate the measurement errors.
-This is particularly relevant, because when we use ZEPE we cannot increase the error due to the measurement because the final state would not be invariant, meaning it cannot be mitigated within this framework.
+Once we have defined the QEP, we can use it as the extrapolation variable in the ZNE method.
+The idea is the same as in standard ZNE: we execute multiple circuits with the same expectation value as the original one but ensuring that the replicas will have larger error probabilities each time.
+To do so, we insert pairs of controlled two-qubit native gates into all connected qubits in the circuit, which in the case of $ibm\_torino$ are controlled-Z gates (CZ).
+We know that a pair of these two-qubit gate is just a unitary operation, so the expectation value remains the same as the original one.
+The addition of this set of gates defines a factor of one.
+We can add as many factors as we want to further amplify the noise, and hence the error probability (see figure 2).
+As we pointed out before, we use the mean QEP to measure the "error", and then we extrapolate the mean QEP to zero to correct our result.
+It is important to note that, as in the standard ZNE, these calculations are compatible with other error mitigation techniques.
+Here, we have used the T-REX mitigation technique Van Den Berg, Minev, and Temme (2022); Chen et al.
+(2019); Bravyi et al.
+(2021) as implemented in Qiskit to eliminate measurement errors.
+This is particularly relevant because when we use ZNE we cannot increase the error due to the measurement because the final state would not be invariant, meaning it cannot be mitigated within this framework.
 Since we have mitigated this source of error by applying T-REX, we did not take it into account while calculating the QEP, assuming it has been largely reduced.
-It has been shown that ZNE performs better when combined with twirling techniques [51].
-Since we aim to compare our method against ZNE, we have also incorporated this error mitigation technique in our calculations.
-Another important issue that we have to take into account is the possible calibration problems explained above.
+ZNE has been shown to perform better when combined with twirling techniques Wallman and Emerson (2016).
+Since we aim to compare our method against standard ZNE, we have also incorporated this error mitigation technique in our calculations.
+Another important issue to take into account is the possible calibration problems explained above.
 Then, the role of warnings identifying if a qubit suffers from calibration issues or unexpected behavior is crucial, since the QEP may be miscalculated, leading to ineffective mitigation.
-Figure 3: Use of ZEPE to mitigate the error in calculating the magnetization of a 68 qubit system after $11$ Trotter steps in $ibm\_torino$.
+Figure 3: Use of QEP to mitigate the error in calculating the magnetization of a 68 qubit system after $11$ Trotter steps with $\delta_{t}=1/4$, $J=\pi$, and $h=0$ in $ibm\_torino$.
 The green points are the raw, factor $1$, factor $2$, factor $3$ values of the magnetization as a function of the mean qubit error probability (QEP).
@@
 In figure 3, we calculate the magnetization of an $11$-Trotter-step Ising system as a function of the average QEP.
-The first green point represents the result of the circuit without ZNE or ZEPE mitigation techniques.
+The first green point represents the result of the circuit without any ZNE mitigation techniques.
 The other three green points correspond to the results obtained after adding a set of two-qubit gates as indicated above (factor $1$, factor $2$, and factor $3$).
@@
 We can see that this method effectively mitigates noise mostly due to the gate errors and qubit state decay.
-It is important to stress that this method, like ZNE, is heuristic, so we cannot guarantee that it will be able to reduce all the noise.
-Figure 4: ZNE and ZEPE magnetizations as a function of the number of Trotter steps.
-Different ZEPE colors correspond to different extrapolation adjustments, i.e.
+It is important to stress that this method, like standard ZNE, is heuristic, so we cannot guarantee that it will be able to reduce all the noise.
+Figure 4: ZNE with QEP magnetizations with $\delta_{t}=1/4$, $J=\pi$, and $h=0$ as a function of the number of Trotter steps.
+Different colors correspond to different extrapolation adjustments, i.e.
 using different set of error amplification factors.
 For instance, blue line labeled as (1,2,3,4,5) corresponds to the linear extrapolation performed with the raw, factor 1, factor 2, factor 3, factor 4, and factor 5 points.
-Finally, we have analyzed ZEPE as a function of the error amplification factor.
+Finally, we have analyzed ZNE as a function of the error amplification factor.
 Different factors can be used to perform the extrapolation.
@@
 Figure 4 shows the results of the extrapolated values using different error amplification factor for the QEP calculation.
-We can see that the best mitigation corresponds to the use of ”raw” and $1$, $2$, and $3$ factors to extrapolate.
+We can see that the best mitigation corresponds to the use of "raw" and $1$, $2$, and $3$ factors to extrapolate.
 Among all the extrapolation options, we have chosen to use the linear extrapolation for two reasons.
@@
 Second, we aim to avoid extrapolations such as the exponential one, which can induce to a significant deviation from the actual expectation value.
-4 The comparison between ZNE and ZEPE
-Once we have developed the ZEPE method, we aim to compare it with the ZNE function implemented in Qiskit.
-As we have previously done, we compute a 68-qubit chain in $ibm\_torino$ quantum system for different numbers of Trotter steps.
-ZEPE has two main advantages in comparison with ZNE.
-First, using ZEPE, we have enhanced control over error and noise, utilizing the quantum computer’s calibration to more precisely monitor noise behavior.
+This is due to the fact that an small error calculating the exponent may induce a huge error in the value we are extrapolating.
+IV Comparison between standard ZNE and ZNE with QEP
+Once we have developed the ZNE with QEP calculation method, we aim to compare it with the standard ZNE function implemented in Qiskit.
+To do so, we will compute the previously explained trotterization for the Ising Hamiltonian for two different cases.
+In the first one, we compute a 68-qubit chain in $ibm\_torino$ quantum system for different numbers of Trotter steps fixing $\delta_{t}=1/4$, $J=\pi$, and $h=0$.
+As we pointed before this correspond to a Clifford point so we can easily know the exact value.
+The second case we will fix $\delta_{t}=1/4$, $J=\pi$ and the number of Trotter steps to $15$, changing the values of $h$.
+This corresponds to non-Clifford points so we will do it for a $32$ qubit system in order to be able to simulate it classically.
+Using QEP has two main advantages in comparison with standard ZNE.
+First, we have enhanced control over error and noise, utilizing the quantum computer’s calibration to more precisely monitor noise behavior.
 Second, increasing the noise is more straightforward and affordable, as it can be achieved by adjusting the circuit depth which enable us to apply it for more complex circuits.
@@
 Left: Absolute value of the total magnetization as a function of the amount of Trotter steps.
-The raw (orange), ZNE (blue), ZEPE (green) and exact value (solid line at $M=68$) are included.
+The raw (orange), standard ZNE (blue), ZNE with QEP (green) and exact value (solid line at $M=68$) are included.
 Right: Relative error of the raw and extrapolated values compaired with the exact theoretical result.
-In figure 5, we can see the absolute magnetization (left) of the raw, ZNE, and ZEPE results, and the absolute error (right) as a function of the number of Trotter steps.
-We can see that our method reduces the error more effectively than ZNE.
-ZEPE performs better for $6$-$16$ Trotter steps.
+Figure 6: Comparison on the reliability between the different error mitigation techniques.
+Absolute value of the total magnetization as a function of the $h$ parameter.
+The raw (orange), standard ZNE (blue), ZNE with QEP (green) and exact value (solid line) are included.
+In figure 5, we can see the absolute magnetization (up) of the raw, i.e.
+without ZNE correction, standard ZNE, and ZNE with QEP results, and the absolute error (bottom) as a function of the number of Trotter steps.
+We can see that our method reduces the error more effectively than standard ZNE.
+It performs better for $6$-$16$ Trotter steps.
 The effectiveness of the method directly depends on the entangle depth of the original quantum circuit.
@@
 On the other hand, if the raw circuit has a mean QEP larger than $0.6$, the error due to circuit time and the error propagation on the gate operations starts to be sufficiently large for an effective mitigation.
-In this regime, ZEPE performs better than ZNE because the latter suffers of incommensurate circuit time execution which yields in a inaccurate manner to account for noise.
-We have shown that ZEPE can be an easily tuneable error mitigation technique.
+In this regime, ZNE with QEP performs better than standard ZNE because the latter suffers of incommensurate circuit time execution which yields in an inaccurate manner to account for noise.
+In figure 6, we present the absolute magnetization for the raw, standard ZNE, and ZNE with QEP results, as a function of the $h$ parameter.
+As in the previous case, the inclusion of QPE improves the effectiveness of standard ZNE.
+We have shown that the use of QPE in ZNE can be an easily tuneable error mitigation technique.
 It requires minimal quantum computational resources and performs better than standard ZNE.
 This new method has been used on IBM quantum computers but can be extended to any type of hardware.
-5 Conclusions
+V Conclusions
 In this work, we have investigated the primary sources of errors in quantum computers to define the QEP, a fundamental metric for quantifying the likelihood of qubit errors.
@@
 The QEP metric provides a systematic approach to assessing errors at the qubit level, offering valuable insights into error propagation and its impact on quantum circuits.
-By leveraging QEP, we aim to enhance the precision and effectiveness of error mitigation strategies in quantum computing.
+By leveraging QEP, we aim to enhance the precision and effectiveness of quantifying the error and mitigate it in quantum computing.
 One of the most critical aspects of the NISQ era is the development of efficient error mitigation techniques to bridge the gap between noisy quantum hardware and the desired quantum advantage.
@@
 However, these approaches have limitations, particularly in their sensitivity to noise models and the complexity of implementation.
-By integrating QEP into this process, we introduced a novel extrapolation method, ZEPE, which refines the standard ZNE approach and enhances its overall performance.
-The results of our study demonstrate that ZEPE offers significant advantages over traditional ZNE methods.
-It requires minimal quantum computational resources while achieving higher accuracy in error mitigation.
-The adaptability and efficiency of ZEPE makes it a promising technique for improving quantum computations in the NISQ era.
-Additionally, ZEPE is easily tunable, allowing users to optimize its parameters based on the specific noise characteristics of a given quantum system.
+By integrating QEP into this process, we improve the standard ZNE approach and enhance its overall performance.
+The results of our study demonstrate that the inclusion of QPE offers significant advantages over traditional ZNE methods.
+As standard ZNE, and in contrast with other mitigation techniques, it requires minimal quantum computational resources while achieving higher accuracy in error mitigation.
+Additionally, it is easily tunable, allowing users to optimize its parameters based on the specific noise characteristics of a given quantum system.
 These attributes make it particularly valuable for practical implementations where computational overhead must be minimized while maintaining high levels of precision.
-Although our research has primarily focused on IBM quantum computers, the ZEPE method is not hardware-specific and can be applied to other quantum computing platforms.
-This flexibility ensures that ZEPE can contribute to a wide range of quantum technologies, enhancing their error resilience and overall performance.
+Although our research has primarily focused on IBM quantum computers, the method is not hardware-specific and can be applied to other quantum computing platforms.
+This flexibility ensures that it can contribute to a wide range of quantum technologies, controlling their errors and enhancing their error resilience and overall performance through the improvement of ZNE.
 Future work could explore further refinements to the technique, including its integration with other error mitigation methods and its application to more complex quantum algorithms.
-By advancing robust and efficient error mitigation strategies, such as ZEPE, we move closer to achieving practical quantum utility, where quantum computations can rival and surpass classical methods in real-world applications.
+By advancing robust and efficient error mitigation strategies, we move closer to achieving practical quantum utility, where quantum computations can rival and surpass classical methods in real-world applications.
 Acknowledgments
@@
 101148213, EATTS).
+Author Declarations
+Conflict of Interest
+The authors have no conflicts to disclose.
+Data Availability
+The data that support the findings of this study are available from the corresponding author upon reasonable request.